Student prepared to start a culture using 2.5*10-3g yeast cells in 1dm3 nutrient broth.

The average mass of yeast cell is 2*10 -11g.

calculate the number of cells in the starter culture 

Give your answer in standard form 


1
Expert's answer
2021-06-07T06:47:07-0400

The number of cells in the starter culture equals:

n(cells) = m(culture) / m(one cell) =  2.5 × 10-3 g / 2 × 10 -11 g = 1.25 × 108


Answer: 108